OpenStudy (anonymous):

how to interpret a real number in exponent i got stuck when doing logarithms : if u take log, for the numbers from 10 to 100, you would get the output [1, 2] which means, \(\log 10 \) to \(\log 100\) will cover all the real numbers range [1, 2]

OpenStudy (anonymous):

should i see the real exponent as, integer + fraction ? and then multiply the integer i get, with a radical ? could somebody plz explain....

OpenStudy (sirm3d):

yes, that's right

OpenStudy (anonymous):

is that the only way

OpenStudy (anonymous):

rational exponent i can split as integer * radical. but what about irrational exponent.. idk if im even ready to think of it. i dont understand the rational exponent itself fully yet.. .

OpenStudy (anonymous):

\(10^{1.6789}\) = ?

OpenStudy (anonymous):

im sure it lies between [1, 2]

OpenStudy (anonymous):

oh sorry the log value of it i mean.. . but how to think of it with out logs.. hmm

OpenStudy (anonymous):

i mean outside logs, where we use this

OpenStudy (sirm3d):

LOG simply strips the base 10 and returns the exponent. so in your case, \[\log_{}10^{1.6789}=1.6789 \]

OpenStudy (anonymous):

yes,

OpenStudy (anonymous):

how to find the value of \(10^{1.6789}\)

OpenStudy (sirm3d):

before the invention of calculators, log is a convenient way of multiplying and or dividing large or small numbers

OpenStudy (anonymous):

ya like, \(10^1 = 10, \) and \(10^2 = 10 \times 10 = 100\) \(10^3 = 10 \times 10 \times 10 = 1000\)

OpenStudy (anonymous):

how can we think of/ or find the value of \(10^{1.6789}\)

OpenStudy (sirm3d):

that is an interesting question if you have to compute it manually

OpenStudy (anonymous):

oh i dont think i wanto compute it manually, i just wanto know how it can be understood, thats all

OpenStudy (anonymous):

\(10^{1.6789} = 10^{1+.6789} = 10^1 + 10^{.6789}\)

OpenStudy (sirm3d):

the equivalent question to that is "what is that number whose logarithm is 1.6789?"

OpenStudy (anonymous):

\(10 + 10^{.6789}\)

OpenStudy (anonymous):

thats right, logarithms are same as exponents

OpenStudy (sirm3d):

the equation is wrong. it should be \[10^{1.6789}=10^{1}10^{0.6789}\]

OpenStudy (anonymous):

so i would like to understand these, either completely in logarithms, or completely in exponetns., without mixing both as we wud go in circles..

OpenStudy (anonymous):

ah thank u yes.s.

OpenStudy (anonymous):

\(10^{1.6789}=10^{1}10^{0.6789} = 10 . 10^{\frac{1}{somthing}}\)

OpenStudy (anonymous):

can i write something like that, and take root of 10

OpenStudy (anonymous):

idk if there is some other elegant way to put these..

OpenStudy (anonymous):

\(10^{1.6789} = 10 . \sqrt[somthing]{10}\)

OpenStudy (anonymous):

and compute the root ? thats how computer does the calculation ?

OpenStudy (anonymous):

\(\huge 10^{1.6789} = 10 . \sqrt[somthing]{10} = 10 . \sqrt[1.5]{10} \)

OpenStudy (anonymous):

i knw i can figure out extracting roots manually.. . it makes bit sense.. but is there any other interpretation for this, easy one ?

OpenStudy (anonymous):

@hartnn @sauravshakya

hartnn (hartnn):

idk, what exactly u wanna do by interpreting, but u can 'guess' that value of 10^1.6789 is around 50, (because its definitely between 10 and 100, and its not near of either of them)